Given that y = x is a solution of (x2 –...


5.

Given that \(y=x\) is a solution of \(\left(x^{2}-x+1\right) y \prime \prime-\left(x^{2}+x\right) y \prime+(x+1) y=0\), a linearly independent solution obtained by reducing the order is given by

\(y=e^{x}(x+1)\)

\(y=e^{x}(x-1)\)

None of them

\(y=x^{2} e^{x}\)

\(y=x e^{x}\)


6.

 If the functions y = x and y = xex are linearly independent solutions of the non-homogeneous second-order linear differential equation with variable coefficients second-order linear differential equation with variable coefficients

\(x^{2} y \prime \prime-x(x+2) y \prime+(x+2) y=x^{3}\), its general solution is given by

None of them

\(y=C_{1}+C_{2} x e^{x}+x^{2}\)

\(y=C_{1} x^{2}+C_{2} x e^{x}-x^{3}\)

\(y=C_{1} x+C_{2} x e^{x}-x^{2}\)

\(y=C_{1} x+C_{2} x^{2} e^{x}-x^{3}\)
Answer #1

option b is true for answer 5 and option a is true for answer 6
Answer #2

(02-271) y - (2+) y +on+ 1) y = 0 - U one solution is y=n. y = ur. .y= untu y = uinta ta un +2m from (n. (? n +1) Cue>+la(w) = nt da -20%-n+1) +7(222-1) 2(2-2+1) 22-1 > ln (w)= nt + ant1 ht) da . > ln(ol =n-2 ln(2) + ln(2²-2+1) + ln (9) In(a) e6) 22 y - 2 (2+2) Y +(2+2) Y =23 yan and y=ner are linearly independent solution. : Yes arteaner .. =nen WCY, 4) = y xlu 1